1. Oxygen gas inside a 1200 litres gas tank has a pressure of 1.0 atm. If the temperature remains constant , how much pressure is needed to reduce its volume by one- half (1/2) ?



1
Expert's answer
2022-04-27T14:23:04-0400

P1 = 1.0 atm

V1 = 1200 L

T1 = T2 = constant

P2 = ???

V2 = (1/2) × V1 = (1/2) × 1200 L = 600 L


Solution:

Since the temperature and amount of gas remain unchanged, Boyle's law can be used.

Boyle's gas law can be expressed as: P1V1 = P2V2

To find the pressure, solve the equation for P2:

P2 = P1V/ V2

P2 = (1.0 atm × 1200 L) / (600 L) = 2.0 atm

P2 = 2.0 atm


Answer: 2.0 atm